OpenStudy (anonymous):

What is the exact value of the expression: sqrt(45)/4 sqrt(81)

OpenStudy (anonymous):

\[\frac{ \sqrt{45} }{ 4\sqrt{81} }=\frac{ \sqrt{9*5} }{ 4(9) }=\frac{ 3\sqrt{5} }{ 36 }=\frac{ \sqrt{5} }{ 12 }\] @vetdoc

OpenStudy (anonymous):

omg thanks